OpenStudy (frostbite):

How can I show that the function f(x)=(1/x)-(cos(x)/sin(x)), ∀x∈R and x≠nπ,n∈Z, is strictly increasing on each interval (nπ,(n+1)π. if the inequality |sin(x)|<|x| is used it have to be proven as well.

OpenStudy (anonymous):

\[f(x)=\frac{1}{x}-\cot x\]???

OpenStudy (anonymous):

so we must prove\[f'(x)=-\frac{1}{x^2}+1+\cot^2 x>0\] for \(n\pi<x<(n+1)\pi\)

OpenStudy (frostbite):

Yes, just posted it in the other form becuase they wrote it that way, but again yes.

OpenStudy (klimenkov):

\(f'(x)=-\frac1{x^2}+\frac1{\sin^2x}\) Using your inequality: \(\frac1{\sin^2x}=\frac1{|\sin x|^2}>\frac1{x^2}\) or \(\frac1{\sin^2x}-\frac1{x^2}>0\) on each interval (nπ,(n+1)π) So, \(f'(x)=-\frac1{x^2}+\frac1{\sin^2x}>0\). It means that your function is strictly increasing on each interval (nπ,(n+1)π).

OpenStudy (frostbite):

How do you get f'(x) to that? Can follow you in the first part, but the second not quite.

OpenStudy (klimenkov):

Can you find the derivative of your function?

OpenStudy (frostbite):

Ok now i got it. Sorry just had to try it by hand.

OpenStudy (klimenkov):

Are you satisfied with the result or you have any questions?

OpenStudy (frostbite):

I think so the reason emm.. can you elaborate the inequality?

OpenStudy (klimenkov):

Yes. What about a geometric proof?

OpenStudy (frostbite):

Would be fine.

OpenStudy (klimenkov):

 There is the circumference or the radius = 1. Area of the sector will be \(\frac x2\) and the area of the triangle will be \(\frac {\sin x} 2\). It is obvious that the area of the sector is bigger than the area of the triangle. So \(\frac{\sin x}2<\frac x 2\) or \(\sin x< x\).

OpenStudy (frostbite):

I see, and the inequality you show in the caculation is just a rewriteing of this inequality?

OpenStudy (klimenkov):

Yes. But I proved it for xє[0,п/2]. You can get your inequality from mine.

OpenStudy (frostbite):

No i ment this one (see the attached file)

OpenStudy (klimenkov):

It is the consequence or the first inequality.

OpenStudy (frostbite):

Ok. I'll say thank you so much. Feel like i got it. Ones again thank you

OpenStudy (klimenkov):

No problem. Ask me again you you'd like!